A 20g weight causes an extension of 0.72 cm in a spring. Calculate the extension caused by a load of 80g weight, assuming Hooke’s law is obeyed.? *
1
Expert's answer
2020-06-29T14:04:09-0400

The Hooke’s law says


"F=kx"

Hence


"\frac{F_2}{F_1}=\frac{x_2}{x_1}"

"x_2=\frac{F_2}{F_1}x_1=\frac{m_2g}{m_1g}x_1=\frac{80}{20}\times 0.72\:\rm cm=2.88 \:cm"